Night Report 2015-03-13

Observer : Morgan Fraser

Support team : Southampton (Mat Smith and others)

TAT : Joe Anderson, Cosimo Inserra

Start of night: Thin cirrus, 2 m/s wind, 65% humidity

Moon from 04.30 onwards at 45% illumination

End of night: Can't see any clouds at dawn, 5 m/s wind, 60% humidity

Seeing: Measured 1" in V at 01.15, subarcsecond for most of night.

Lost ~1hr total due to technical issue with rotator (not caused by going to zenith - this is

an intermittent fault which has started happening over the last couple of months).

Afternoon calibrations: SOFI arcs and specflats

23.10 Sunset

23.25 HIP 016853 (Telluric standard for PSNJ07361576-6930230 )

BG+RG

23.30 PSNJ07361576-6930230

SOFI JHK imaging, BG+RG spectra

00.23 End twilight

00.55 Switch to EFOSC

01.04 GD71 (standard)

01.21 Telescope error - system restated

01.29 OGLE-2015-SN-043

BVRi imaging, seeing ~1" at start of sequence

02.40 OGLE-2015-SN-035

Gr#13 2x1800s

Telescope error after first OB (03:17). Problem is connected to rotator, ESO are

aware of this as an intermittent issue, but do not know the cause yet.

04.10 PSNJ07361576-6930230

600s Gr#11,16

Calibration unit for internal flats did not move, TO had to insert manually.

04.49 PS15br

2x1500s Gr#16

05.56 LSQ15kp

2x1800s Gr#11 (taking Gr#16 later when target has passed zenith to avoid rotator limit)

07.06 - Second OB ended early after ~1700s due to telescope error (this was not due to

being too close to zenith - elevation was <80 deg when problem occured).

07.12 PSNJ14095513+1731556

2x1800s Gr#13

08.18 LSQ15kp

2x1800s Gr#16

09.21 Twilight

09.34 EG274 (standard)

09.47 LSQ15so

Gr#13 classification, Ia, z=0.07, a few days post max.

10.05 LSQ15st

Gr#13 classification, only got 300s as sky was getting brighter.

Ia, z=0.06, post max (best fitting templates are pre-max, but LC is clearly post-peak)

10.34 Sunrise (too bright for skyflats)